# Multi-Scale Attention and Gated Shifting for Fine-Grained Event Spotting in Videos

### v1

@@ 5 Discussion
The design of MSAGSM enables rich temporal modeling with minimal additional computational overhead, making it a practical choice for real-world applications. We observe that MSAGSM performs particularly well when paired with both stronger backbone architectures and simpler models, suggesting that stronger feature representations further amplify the module’s ability to capture temporal dynamics.
In addition, we introduce the TTA dataset—the first table tennis dataset for PES—which brings unique challenges due to its high event density and rapid scene changes. We hope this dataset will serve as a new benchmark to advance research in this domain.
@@
While our work focuses on PES, MSAGSM has potential for broader video understanding tasks, such as video action recognition (VAR). Unlike PES, VAR typically uses sparse frame sampling, which may limit its temporal precision. Future work will explore how fine-grained modules like MSAGSM can bridge this gap and enhance VAR performance.
Overall, we envision MSAGSM as a lightweight, general-purpose temporal module that can be easily integrated into existing pipelines to improve fine-grained temporal reasoning in video understanding.
@@ 6 Conclusion
We introduce a simple yet effective temporal module, the MSAGSM, which can be seamlessly integrated into any existing 2D backbone feature extractor to enable multi-scale temporal modeling. MSAGSM leverages both multi-scale temporal shifts and an attention mechanism to selectively aggregate information across frames at varying temporal distances. We conduct extensive evaluations across five fine-grained sports video datasets for PES, demonstrating that MSAGSM consistently enhances model performance and achieves state-of-the-art results. For instance, when integrated into E2E-Spot, MSAGSM yields an absolute improvement of +3.08% mAP@1 on the TTA dataset, while introducing only 0.01M additional parameters. Furthermore, we propose TTA, a new PES benchmark dataset in the domain of table tennis, designed to encourage future research on fine-grained and high-frequency sports video understanding.
Appendix A Implementation Details
A.1 Backbone Model
@@ A.2 Baselines
A.3 Training Configuration
All experiments were conducted on the same machine equipped with L40S GPUs to ensure consistency. The overall training configuration follows E2E-Spot: each epoch randomly samples 5,000 clips to reduce training time while maintaining good coverage of the dataset. To mitigate the imbalance between event and non-event frames, we applied a weight of 5× to the event classes.
We trained for 50 epochs using AdamW with an initial learning rate of $1\times 10^{-3}$, a batch size of 8, and 3 warm-up epochs. Cross-entropy loss was used, with input tensors shaped as $[B,T,C,H,W]$ (batch, time, channels, height, width) and output tensors as $[B,T,C]$, representing the frame-wise class probabilities after softmax.
A.4 Data Augmentations
@@
Table 3: Performance comparison (mAP %) under different temporal tolerances ($\delta=1$ and $\delta=2$) for T-DEED and UGLF under ResNet-18 backbone. Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp Module TTA Tennis FS_Comp FS_Perf FineDiving Parameters
$\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2
E2E-Spot ResNet-18 GSM 69.13 78.37 90.43 96.72 77.31 87.52 82.75 94.56 64.63 84.79 14.35M
E2E-Spot ResNet-18 GSF 66.42 80.75 90.48 97.01 73.21 87.37 78.50 90.58 64.25 84.01 14.35M
E2E-Spot ResNet-18 Ours 70.49 79.38 90.81 97.52 78.29 90.93 85.23 96.62 67.17 84.95 14.37M
T-DEED ResNet-18 GSM 66.78 76.13 89.46 97.19 76.65 87.02 83.67 95.35 63.11 83.34 37.34M
T-DEED ResNet-18 GSF 68.84 75.59 88.48 97.22 75.58 88.87 83.09 96.22 68.68 85.21 37.34M
T-DEED ResNet-18 Ours 69.92 77.33 87.18 95.85 76.17 88.57 82.58 96.35 66.78 84.88 37.37M
UGLF ResNet-18 GSM 60.05 69.50 89.38 97.14 75.93 90.16 81.28 92.60 65.88 84.92 14.35M
UGLF ResNet-18 GSF 62.58 68.57 89.11 97.31 74.88 88.49 82.34 94.27 65.56 84.56 14.35M
UGLF ResNet-18 Ours 62.30 72.14 91.06 97.28 77.21 90.25 81.46 94.84 68.98 87.60 14.37M
Appendix B TTA Dataset
@@
All videos in the TTA dataset were recorded and provided by the Paralympics Table Tennis Australia team. The dataset is intended strictly for research purposes and will be made publicly available upon request and approval.
@@
Table 4: Performance comparison (mAP %) under strict temporal tolerance ($\delta=0$) for T-DEED and UGLF under RegNetY-200 backbone. Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp Module TTA Tennis FS_Comp FS_Perf FineDiving Parameters
E2E-Spot RegNetY-200 GSM 34.36 51.42 36.44 39.32 27.31 4.45M
E2E-Spot RegNetY-200 GSF 36.17 49.68 36.25 40.07 29.41 4.45M
E2E-Spot RegNetY-200 Ours 36.43 55.51 38.34 46.54 29.83 4.50M
T-DEED RegNetY-200 GSM 33.32 48.82 38.61 26.86 26.86 16.42M
T-DEED RegNetY-200 GSF 26.27 50.08 38.83 38.89 28.50 16.42M
T-DEED RegNetY-200 Ours 28.06 52.76 39.06 44.66 26.88 16.48M
UGLF RegNetY-200 GSM 34.36 50.31 36.99 38.29 27.11 4.45M
UGLF RegNetY-200 GSF 36.17 50.94 34.84 37.41 29.81 4.45M
UGLF RegNetY-200 Ours 33.93 53.10 39.66 46.59 29.59 4.50M
B.1 Dataset Annotation Details
@@
Table 5: Performance comparison (mAP %) under strict temporal tolerance ($\delta=0$) for T-DEED and UGLF under ResNet-18 backbone. Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp Module TTA Tennis FS_Comp FS_Perf FineDiving Parameters
E2E-Spot ResNet-18 GSM 40.30 49.74 37.80 43.73 26.53 14.35M
E2E-Spot ResNet-18 GSF 39.11 49.32 35.50 39.04 24.31 14.35M
E2E-Spot ResNet-18 Ours 43.81 52.82 40.49 45.75 30.43 14.37M
T-DEED ResNet-18 GSM 39.81 48.09 38.64 42.82 25.86 37.34M
T-DEED ResNet-18 GSF 38.26 45.57 35.54 42.03 28.81 37.34M
T-DEED ResNet-18 Ours 39.51 51.56 36.89 40.90 23.68 37.37M
UGLF ResNet-18 GSM 32.89 49.05 36.57 42.30 24.53 14.35M
UGLF ResNet-18 GSF 36.98 47.59 34.22 41.73 22.42 14.35M
UGLF ResNet-18 Ours 37.14 52.14 40.69 40.55 29.56 14.37M
Appendix C Additional Experiments
@@
C.1 Different Backbone
@@
C.2 Exact Frame Prediction
In PES settings, mAP@0 measures whether the prediction lands exactly on the ground-truth frame. However, this is often ambiguous even for human annotators, so mAP@0 is considered a supplementary result. These results are presented in Table 4 for RegNetY-200 and Table 5 for ResNet-18.
We observe that even under the strict mAP@0 setting, MSAGSM achieves strong results: on RegNetY-200, it achieves the best result in 12 out of 15 cases, and on ResNet-18, in 10 out of 15 cases. This suggests that the module is particularly effective at pinpointing exact event moments when integrated into simpler architectures such as E2E-Spot, as it can fully leverage its multi-scale shifts and spatial attention. In contrast, in more complex models like T-DEED—which already incorporate advanced temporal mechanisms—MSAGSM may partially overlap or conflict with existing components, slightly reducing its relative effectiveness.
@@ C.2 Exact Frame Prediction
Setting Value TTA FineDiving Params
mAP@1 mAP@2 mAP@1 mAP@2
@@
Heads 1 60.99 75.38 66.36 83.47 4.50M
Heads 2 69.50 (+8.51) 78.12 (+2.74) 71.03 (+4.67) 87.26 (+3.79) 4.51M
Heads 3 65.87 (+4.88) 75.87 (+0.49) 67.93 (+1.57) 85.11 (+1.64) 4.51M
Heads 4 61.01 (+0.02) 72.98 (–2.40) 68.72 (+2.36) 85.16 (+1.69) 4.52M
Heads 8 64.19 (+3.20) 72.94 (–2.44) 70.06 (+3.70) 85.91 (+2.44) 4.55M
@@
Dilations [1,2] 60.78 71.10 68.56 85.49 4.48M
Distance [1,2,3] 69.50 (+8.72) 78.12 (+7.02) 71.03 (+2.47) 87.26 (+1.77) 4.51M
Dilations [1,3,5] 68.14 (+7.36) 75.26 (+4.16) 68.18 (-0.34) 85.26 (–0.23) 4.51M
Distance [1,…,4] 64.18 (+3.40) 71.59 (+0.49) 69.02 (+0.46) 85.17 (-0.32) 4.53M
Dilations [1,2,4,6] 68.27 (+7.49) 75.07 (+3.97) 69.85 (+1.29) 87.02 (+1.53) 4.53M
Distance [1,…,5] 67.54 (+6.76) 67.41 (-3.69) 69.26 (+0.70) 85.26 (-0.23) 4.55M
Distance [1,…,6] 64.03 (+3.25) 75.38 (+4.28) 69.84 (+1.28) 85.34 (-0.15) 4.57M
Dilations [1,…,9] 67.32 (+6.54) 76.44 (+5.34) 65.47 (–3.09) 82.80 (–3.97) 4.63M
C.3 Different Temporal Window Setting
Here, we present a more detailed evaluation of temporal window settings for MSAGSM, including wider ranges such as $[1,\dots,9]$ and skipped patterns like $[1,2,4,6]$. As longer temporal distances are incorporated, the improvements become unstable and often diminish, suggesting that shifting information from nearby frames is sufficient in most cases. Notably, the configuration $[1,2,3]$ consistently performs best, effectively balancing short-term and moderate long-term dependencies. Smaller windows (like $[1,2]$) fail to capture enough context, while larger windows (like $[1,\dots,9]$) include redundant or noisy temporal information, which can confuse the model and harm precision. Thus, moderate dilation allows the model to attend to meaningful temporal cues without overfitting to irrelevant frames.
C.4 Class-wise Spotting Performance

### v2

@@ 5 Limitations and Future Work
MSAGSM enables rich temporal modeling with minimal overhead, making it a practical choice for real-world applications. We observe that it performs particularly well with both strong backbones and lightweight models, suggesting that better feature representations further amplify its effectiveness.
@@
Although our work focuses on PES, MSAGSM could also benefit broader video understanding tasks such as action recognition, where sparse frame sampling limits temporal precision. Future work may explore how fine-grained modules like MSAGSM can bridge this gap.
We acknowledge two key limitations. First, MSAGSM is sensitive to hyperparameters (e.g., temporal distance, attention heads), requiring dataset-specific tuning and reducing plug-and-play generalization across datasets. Second, per-class precision–recall analysis on TTA (Supp. §3.6) shows frequent confusion between forehand and backhand strokes due to their visual similarity and player handedness, whereas distinct events like bounces are recognized more reliably.
Future directions include adaptive strategies for automatically learning temporal distances and methods to better handle fine-grained ambiguities and viewpoint variations.
@@ 6 Conclusion
We propose MSAGSM, a lightweight temporal module for 2D backbones that combines temporal shifts and attention. Across five fine-grained PES datasets, it consistently improves performance, boosting mAP@0 by +4.09 and mAP@1 by +3.08% on TTA with only 0.01M extra parameters. We also present TTA, a new table tennis PES benchmark with realistic recordings and high event density, to support future research in fine-grained sports video understanding.
Appendix A Notation Table
For clarity, we provide a summary of the main symbols used throughout the paper in Table 4. This table lists the definitions of key variables (scalars, vectors, matrices, and tensors) along with evaluation parameters such as class weights and temporal tolerance, to ensure consistent interpretation of our mathematical notation.
Table 4: Summary of notation used in the paper.
Symbol Description
$t$ Frame index
$L$ Clip length (number of frames per input)
$x_{t}$ Input frame at time $t$
$F_{t}\in\mathbb{R}^{C\times H\times W}$ Feature map at frame $t$ ($C$ channels, $H\times W$ spatial size)
$C,H,W$ Number of channels, height, and width of feature maps
$X_{\text{left}},X_{\text{right}}$ Partitioned input features for left and right shifts
$G_{\text{left}},G_{\text{right}}$ Learned gating maps (via 3D convolution + $\tanh$)
$\odot$ Element-wise (Hadamard) product
$\operatorname{Roll}(X,\pm\delta)$ Circular shift (temporal roll) of features by $\delta$ frames left/right
$\delta$ Temporal shift distance (in frames)
$\operatorname{Concat}(\cdot,\cdot)$ Concatenation of feature groups along the channel dimension
$\hat{X}$ Updated features after gated shifts and concatenation
$\hat{X}^{(i)}$ Output from the $i$-th temporal scale
$w_{i}$ Learnable weight for scale $i$ ($w_{i}\geq 0$, $\sum_{i}w_{i}=1$)
$\hat{X}_{\text{final}}$ Weighted sum of multi-scale shifted features
$y_{t},\hat{y}_{t}$ Ground-truth event label and predicted score at frame $t$
$w$ Class weighting factor in cross-entropy loss
$\delta$ (eval) Temporal tolerance used in evaluation metrics
Appendix B Implementation Details
B.1 Backbone Model
@@ B.2 Baselines
We included ASFormer [50] as a transformer baseline, following prior work [12, 44, 27]. In earlier studies [12], ASFormer was evaluated using TSP as the feature extractor, pretrained on video action recognition datasets and then finetuned on PES. To ensure fairness, we instead used RegNetY as the feature extractor for both ASFormer and ASTRA [43], since ASTRA was originally designed to operate on pre-extracted SoccerNet features, including audio inputs.
@@
B.3 Training Configuration
All experiments were conducted on the same machine with a single L40S GPU. For fairness, we closely follow the training protocol of E2E-Spot [12]. Each epoch samples 5,000 clips to reduce training time while maintaining coverage of the dataset, and event frames are weighted 5× in the loss to address class imbalance. We train for 50 epochs using AdamW with an initial learning rate of $1\times 10^{-3}$, a batch size of 8, and 3 warm-up epochs followed by cosine decay. As in E2E-Spot, cross-entropy loss is applied, with inputs shaped $[B,T,C,H,W]$ (batch, time, channels, height, width) and outputs $[B,T,C]$, representing frame-wise class probabilities after softmax.
B.4 Data Augmentations
@@
Table 5: Performance comparison (mAP %) under different temporal tolerances ($\delta=0,1,2$) for T-DEED and UGLF under ResNet-18 backbone. Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp TTA Tennis FS_Comp FS_Perf FineDiving P (M) / GFLOPs
$\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2
E2E-Spot ResNet-18 GSM 40.30 69.13 78.37 49.74 90.43 96.72 37.80 77.31 87.52 43.73 82.75 94.56 26.53 64.63 84.79 14.35 / 367.04
E2E-Spot ResNet-18 GSF 39.11 66.42 80.75 49.32 90.48 97.01 35.50 73.21 87.37 39.04 78.50 90.58 24.31 64.25 84.01 14.35 / 367.04
E2E-Spot ResNet-18 Ours 43.81 70.49 79.38 52.82 90.81 97.52 40.49 78.29 90.93 45.75 85.23 96.62 30.43 67.17 84.95 14.37 / 370.40
T-DEED ResNet-18 GSM 39.81 66.78 76.13 48.09 89.46 97.19 38.64 76.65 87.02 42.82 83.67 95.35 25.86 63.11 83.34 37.34 / 369.26
T-DEED ResNet-18 GSF 38.26 68.84 75.59 45.57 88.48 97.22 35.54 75.58 88.87 42.03 83.09 96.22 28.81 68.68 85.21 37.34 / 369.26
T-DEED ResNet-18 Ours 39.51 69.92 77.33 51.56 87.18 95.85 36.89 76.17 88.57 40.90 82.58 96.35 23.68 66.78 84.88 37.37 / 372.62
UGLF ResNet-18 GSM 32.89 60.05 69.50 49.05 89.38 97.14 36.57 75.93 90.16 42.30 81.28 92.60 24.53 65.88 84.92 14.35 / 367.04
UGLF ResNet-18 GSF 36.98 62.58 68.57 47.59 89.11 97.31 34.22 74.88 88.49 41.73 82.34 94.27 22.42 65.56 84.56 14.35 /367.04
UGLF ResNet-18 Ours 37.14 62.30 72.14 52.14 91.06 97.28 40.69 77.21 90.25 40.55 81.46 94.84 29.56 68.98 87.60 14.37 / 370.40
Appendix C TTA Dataset
@@
All videos in the TTA dataset were recorded and provided by the Paralympics Table Tennis Australia team. The dataset is intended strictly for research purposes and will be available upon request and approval.
@@
C.1 Dataset Annotation Details
@@
Appendix D Additional Experiments
@@
Table 6: Ablation study on attention head count and dilation range using the MSAGSM module on the TTA and FineDiving datasets in the Validation Set. All results are reported in mAP (%). We also report the average performance across the two datasets.
Setting Value TTA FineDiving Average P (M) / GFLOPs
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2 Avg@0 Avg@1 Avg@2
Varying Number of Attention Heads (Dilations = [1,2,3])
Heads 1 41.97 68.08 72.26 31.72 73.03 88.28 36.85 70.56 80.27 4.50 / 44.68
Heads 2 38.78 71.67 79.37 33.72 73.58 89.77 36.25 72.63 84.57 4.51 / 45.06
Heads 3 38.48 67.23 71.59 30.68 71.29 86.32 34.58 69.26 78.96 4.51 / 45.42
Heads 4 40.13 67.20 70.87 32.98 74.81 87.81 36.56 71.01 79.34 4.52 / 45.78
Heads 8 39.91 66.61 73.54 30.26 72.67 88.20 35.09 69.64 80.87 4.55 / 47.24
Varying Temporal Window Size (Heads = 2)
Distance [1,2] 34.07 64.46 70.68 31.63 74.08 87.63 32.85 69.27 79.16 4.48 / 43.98
Distance [1,2,3] 38.78 71.67 79.37 33.72 73.58 89.77 36.25 72.63 84.57 4.51 / 45.06
Distance [1,…,4] 37.54 65.48 71.84 30.95 72.72 86.47 34.25 69.10 79.16 4.53 / 46.12
Distance [1,…,5] 42.45 69.47 74.16 31.77 74.19 88.41 37.11 71.83 81.29 4.55 / 47.18
Distance [1,…,6] 36.19 67.87 74.91 32.37 74.11 87.79 34.28 70.99 81.35 4.57 / 48.26
Distance [1,…,9] 45.20 74.37 79.10 32.78 70.21 85.37 38.99 72.29 82.24 4.63 / 51.46
D.1 Different Backbone
@@
D.2 Exact Frame Prediction
In PES settings, mAP@0 measures whether the prediction lands exactly on the ground-truth frame. However, this is often ambiguous even for human annotators, so mAP@0 is considered a supplementary result. These results are presented in Table 5 for ResNet-18.
We observe that even under the strict mAP@0 setting, MSAGSM achieves strong results: on ResNet-18, in 10 out of 15 cases. This suggests that the module is particularly effective at pinpointing exact event moments when integrated into simpler architectures such as E2E-Spot, as it can fully leverage its multi-scale shifts and spatial attention. In contrast, in more complex models like T-DEED—which already incorporate advanced temporal mechanisms—MSAGSM may partially overlap or conflict with existing components, slightly reducing its relative effectiveness.
Table 7: Sensitivity analysis of the cross-entropy (CE) loss weighting factor on the TTA and FineDiving datasets. All results are reported in mAP (%).
Loss / Weight TTA FineDiving
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2
CE (1) 29.65 58.42 69.56 31.61 65.13 85.75
CE (3) 32.04 60.52 73.38 27.09 67.08 83.16
CE (5) 36.43 69.50 78.12 29.83 71.03 87.26
CE (7) 35.50 67.83 77.15 29.76 68.97 86.63
CE (9) 30.76 60.09 73.40 32.77 69.43 86.61
Focal 32.54 58.65 71.72 30.71 67.93 85.94
@@ D.2 Exact Frame Prediction
Setting Value TTA FineDiving P (M) / GFLOPs
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2
@@
Heads 1 32.90 60.99 75.38 29.91 66.36 83.47 4.50 / 44.68
Heads 2 36.43 69.50 78.12 29.83 71.03 87.26 4.51 / 45.06
Heads 3 35.98 65.87 75.87 28.68 67.93 85.11 4.51 / 45.42
Heads 4 30.34 61.01 72.98 30.19 68.72 85.16 4.52 / 45.78
Heads 8 35.26 64.19 72.94 29.03 70.06 85.91 4.55 / 47.24
@@
Distance [1,2] 29.64 60.78 71.10 30.18 68.56 85.49 4.48 / 43.98
Distance [1,2,3] 36.43 69.50 78.12 29.83 71.03 87.26 4.51 / 45.06
Distance [1,3,5] 29.83 68.14 75.26 29.69 68.18 85.26 4.51 / 45.06
Distance [1,…,4] 30.06 64.18 71.59 28.85 69.02 85.17 4.53 / 46.12
Distance [1,2,4,6] 39.77 68.27 75.07 32.68 69.85 87.02 4.53 / 46.12
Distance [1,…,5] 29.41 67.54 67.41 30.20 69.26 85.26 4.55 / 47.18
Distance [1,…,6] 35.07 64.03 75.38 29.55 69.84 85.34 4.57 / 48.26
Distance [1,…,9] 41.02 67.32 76.44 27.76 65.47 82.80 4.63 / 51.46
D.3 Validation-Based Hyperparameter Selection
We selected our default hyperparameters by running experiments on validation sets from two datasets. As shown in Table 6, while the optimal configuration may vary across datasets due to differing characteristics, the default setting of $[1,2,3]$ with two attention heads provides the most consistent performance with minimal complexity overhead.
D.4 Loss Function Sensitivity Analysis
We conducted a sensitivity analysis of the Cross-Entropy (CE) loss weighting for the foreground (event) class, and also compared CE with focal loss on two datasets (Table 7). We observe that a weighting factor of 5 yields the best performance on both datasets, outperforming focal loss. This aligns with the standard practice reported in prior work such as E2E-Spot [12], T-DEED [44], and UGLF [37].
D.5 Different Temporal Window Setting
Here, we present a more detailed evaluation of temporal window settings for MSAGSM as shown in Table 8, including wider ranges such as $1,\dots,9]$ and skipped patterns like $[1,2,4,6]$. As longer temporal distances are incorporated, the improvements become unstable and often diminish, suggesting that shifting information from nearby frames is sufficient in most cases. Notably, the configuration $[1,2,3]$ consistently performs best, effectively balancing short-term and moderate long-term dependencies. Smaller windows (like $[1,2]$) fail to capture enough context, while larger windows (like $1,\dots,9]$) include redundant or noisy temporal information, which can confuse the model and harm precision. Thus, moderate dilation allows the model to attend to meaningful temporal cues without overfitting to irrelevant frames.
D.6 Class-wise Spotting Performance
@@
D.7 Visualization of the Model
We visualize attention maps using Grad-CAM to highlight how each model attends to critical regions during precise event spotting as shown in Figure 9. For table-bounce events under occlusion, baselines such as E2E, T-DEED, and UGLF often fail to focus on the relevant regions, whereas our approach consistently attends to the table area, demonstrating its ability to leverage long-range temporal context and effective attention mechanisms. Similarly, for jump-landing events, our model correctly emphasizes the skater’s feet, rather than irrelevant regions.
Figure 9: Grad-CAM visualizations comparing E2E-MSAGSM with baseline models (E2E, T-DEED, and UGLF) across different event types. For close table bounces under occlusion, only our model consistently attends to the correct table region, while baselines often focus elsewhere. For far table bounces, E2E-MSAGSM maintains attention on the bounce area, whereas others drift toward irrelevant regions. For jump-landing events, our model emphasizes the skater’s feet, in contrast to baselines that spread attention across less informative areas. These visualizations demonstrate the effectiveness of our module in capturing both long-range temporal context and spatially discriminative cues.